OpenStudy (anonymous):

Find the four arithmetic means between 12 and -13. A. 6, -1, -3, -8 B. 6, 2, -3, -8 C. 7, -1, -3, -8 D. 7, 2, -3, -8

OpenStudy (anonymous):

\[a _{1} ,a _{2} ,a _{3} ,a _{4}~be ~four~arithmetic~means~between~12,-13\]

OpenStudy (anonymous):

then \[12,a _{1} ,a _{2} ,a _{3} ,a _{4} ,-13 ~are~in~A.P.\] a=12 l=-13 no. of terms n=6 let d be c.d. l=a+(n-1)d -13=12+(6-1)d -25=5d \[d=\frac{ -25 }{ 5 }=-5\] \[a _{1}=12+d=12-5=7,a _{2}=7+d=7-5=2,\] \[a _{3}=2+d=2-5=-3\] \[a _{4}=-3+d=-3-5=-8\]